OpenStudy (anonymous):

Help: 1. What is the slope of the line passing through the points (2, 7) and (–1, 3)? (1 point) a.2/7 b.3/4 c.4/3 d.1/3

OpenStudy (hoblos):

slope = (3-7)/(-1-2) = -4/(-3) = 4/3

OpenStudy (anonymous):